Ask
your own question, for FREE!
Mathematics
8 Online
You have a 4-card deck containing a queen, a jack, a 10, and a 9. You draw a random card, then put it back and draw a second random card. Use a tree diagram to calculate the probability that you draw exactly 1 queen.
Still Need Help?
Join the QuestionCove community and study together with friends!
p(something else and then queen) = 3/4 * 1/4 = 3/16 = p(queen than something else) p=3/8
hope this is easier then before lol
Yes thank you so much! :)
welcome
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 hour ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal